Can the Solow model be improved by explicitly recognising a role for human and physical capital? Really struggling with how to write a conclusion for this.
Answer #1

The slow model is an economic model of long-run economic growth under neoclassical economics.

There are two types of capital,physical &human capital.

Physical capital k(t+1)=k(t)=sy(t)

Human capital h(t+1)-h(t)=ay(t)

Think of qy(t) as the quantity of physical resources spent on education and training.

Solow's source of economic growth estimates of separate effects on technological change,capital and labor .

Lags in the diffusion of return of knowledge. Differences in real income might shrink as poor countries receives better technology and infrastructure.